OpenStudy (anonymous):

Mandy was laying tile in her kitchen. She needed 2 and three-fourths rows with 3 and one-half tiles per row to cover the entire floor. How many total tiles did she need? Multiply to find out. PLEASE HELP ME PLEASE!!!!!

OpenStudy (blurbendy):

2 3/4 * 3 1/2 = 11/4 * 7/2 = 77/8 = 9 and 5/8 tiles